in the complex with formula Mcl3.4H2O, the coordination number of the metal M is six and there is no molecule of hydration in it . the volume of 0.1 M AgNo3 solution needed to precipitate the free chloride ions in 200ml of 0.01 M solution of the complex is ?


Grade:11

1 Answers

ruchi yadav
10 years ago

Since the coordination no. is 6 and there is no water of hydration, the actual formula is [MCl2(H2O)]Cl.

Thus , no. of ionisable Cl-in 1 mol of complex is 1.

Total no of moles of the given compound= (200/1000)*0.01 =2*10-3= no. of moles of chloride= no. of moles of AgNO3required.

Thus, volume of AgNO3required= 2*10-3/0.1 = 2*10-2L =20 mL.


Thank You

Ruchi

Askiitians Faculty